OpenStudy (anonymous):

Which equation matches the statement “The product of 3 and the difference of a number and 10 is 15?"

OpenStudy (anonymous):

what are the choices

OpenStudy (anonymous):

3 + (x – 10) = 15 3x – 10 = 15 3 (x – 10) = 15 3three x over ten = 15

OpenStudy (anonymous):

2. –6q + 5.8 = 23.8 (1 point) q = 3 q = –3 q = –1.83 q = 4.93

OpenStudy (anonymous):

-6q+5.8=23.8 -5.8 -5.8 -6q=18 divide both sides by -6 q=-3

OpenStudy (anonymous):

and I believe the other one is 3(x-10)=15